### v1

# Leveraging Core and Uncore Frequency Scaling for Power-Efficient Serverless Workflows

Abstract.
Serverless workflows have emerged in FaaS platforms to represent the operational structure of traditional applications. With latency propagation effects becoming increasingly prominent, step-wise resource tuning is required to address the end-to-end Quality-of-Service (QoS) requirements. Modern processors’ allowance for fine-grained Dynamic Voltage and Frequency Scaling (DVFS), coupled with the intermittent nature of serverless workflows presents a unique opportunity to reduce power while meeting QoS.
In this paper, we introduce a QoS-aware DVFS framework for serverless workflows. $\Omega$kypous regulates the end-to-end latency of serverless workflows by supplying to the system the Core/Uncore frequency combination that minimizes power consumption. With Uncore DVFS enriching the efficient power configurations space, we devise a grey-box model that accurately projects functions’ execution latency and power, to the applied Core and Uncore frequency combination. To the best of our knowledge, $\Omega$kypous is the first work that leverages Core and Uncore DVFS as an integral part of serverless workflows. Our evaluation on the analyzed Azure Trace, against state-of-the-art (SotA) power managers, demonstrates an average power consumption reduction of 9% (up to 21%) while minimizing QoS violations.
1. Introduction
Serverless computing platforms (e.g., AWS Lambda (awslambda, ), Microsoft Azure Functions (azurefunctions, ), etc.) continue to grow in popularity, as they alleviate the complexities of infrastructure management and resource provisioning. Serverless relies on the execution of short-lived tasks (functions), which are orchestrated into dependency graph workflows (chains), thus, enabling the construction of complex applications through interconnected functions. These workflows are managed by backend services (e.g., AWS Step Functions (awsstep, ), Alibaba Serverless Workflow (cloudflow, ), Azure Durable Functions (azure-durable, ), OpenWhisk (openwhisk-site, )) which handle the triggering, execution sequencing, and resource management of each function.
The strategy of decomposing applications into short functions forms a key enabler for more fine-grained resource management policies, tailored to the unique characteristics of each function. As applications experience varying phases throughout their lifecycle  (masouros2020rusty, ; kannan2019caliper, ), the intermittent nature of serverless workflows allows for “right-sizing” resources for each function, aligning with different optimization goals. While prior research has explored this for Quality-of-Service (QoS) guarantees (mahgoub2021sonic, ; llama, ; zhou2022aquatope, ; moghimi2023parrotfish, ; tariq2020sequoia, ), there is minimal focus on serverless computing’s potential in managing power consumption. Notably, power consumption is increasingly significant for data center facilities providers  (lo2014towards, ; kontorinis2012managing, ; haque2017exploiting, ), with operational expenses closely tied to the power needs of these facilities. For instance, Google incorporates CPU speed into their billing policies  (wang2018peeking, ), and IBM has recently open-sourced a cloud-native power monitor (kepler, ; amaral2023kepler, ), indicating a growing awareness and need for efficient power management in serverless environments.
Power management in the cloud: From a system level perspective, power management in cloud infrastructures is
Figure 1. UFS expands the Latency-Power Pareto front
commonly tackled through the implementation of Dynamic Voltage and Frequency Scaling (DVFS) techniques (retail, ; gemini2020zhou, ; kasture2014ubik, ), the utilization of Asymmetric Multicore Processors (AMPs) (haque2017exploiting, ) and the integration of sleep state management  (chou2019mudpm, ; sharafzadeh2019yawn, ; agilewatts, ). Numerous research efforts (retail, ; gemini2020zhou, ; kasture2014ubik, ) focus on optimizing power management without compromising QoS constraints, yet they typically target either single-stage (microservice-level) or monolithic deployments. Such approaches do not adequately account for second-level effects, intrinsic to serverless architectures, such as function-level timing slacks arising either from the execution of the function itself or from the latency propagation between interdependent functions (tzenetopoulos2023dvfaas, ; seqclock, ). Moreover, these models predict power use and latency from the core’s clock frequency, representing just about a third of a server’s total power consumption (meisner2009powernap, ).
However, with post-Haswell era Intel® processors offering independent frequency control for both the CPU core (CFS) and Uncore (UFS) components, there is a novel opportunity to further enhance infrastructure efficiency, through the coordinated configuration of CFS and UFS, allowing for a more refined balance between performance and power consumption. Figure 1 shows the power - latency profile for a typical serverless function (thumbnailer (sebs, )) when operating under differing CFS and UFS configurations. As shown, the coordinated selection of CFS & UFS reveals an expanded solution space of higher quality that i) completely dominates the conventional CFS solution, and ii) enables iso-latency configurations with up to 25% power gains.
Harnessing the “power” of serverless functions: The inherent flexibility and fine-granularity of serverless computing allows for a “just-right” approach, where resources are optimized to exactly meet QoS objectives, i.e., being QoS-precise. Coupled with the extended power/latency surface offered by CFS/UFS configurations, this strategy presents a significant opportunity for power savings through the careful adjustment of operating clock frequencies based on real-time performance metrics and workload characteristics. However, realizing this potential is not without challenges, as it necessitates sophisticated modeling to understand the performance and power impact of each setting at the function level (§3). The diverse resource requirements of serverless functions, along with their varied input workloads (sahraei2023xfaas, ; shahrad2019architectural, ; romero2021faa, ) make predicting latency and power consumption extremely challenging. Existing approaches (gemini2020zhou, ; zhou2022aquatope, ; retail, ) attempt to leverage latency estimation for QoS-constrained power management, but they face practical limitations. Some require detailed per-function modeling (retail, ), while others assume a linear relationship between latency and frequency (gemini2020zhou, ), or require multiple iterations for configuration updates (zhou2022aquatope, ). Notably, according to Azure’s insights (zhang2021faster, ), serverless workflows’ invocation count may not be adequate to effectively train ML-based latency estimation models in a per-configuration manner. Hence, there is a need for practical, more mechanistic, less data-driven, and universally applicable approaches to address resource management requirements promptly.
Our work. In this paper, we present $\Omega$kypous, a novel, QoS-precise function-granular DVFS control framework that leverages the intermittent nature of serverless workflows to minimize power consumption while guaranteeing the end-to-end QoS requirement. We employ an adaptive frequency selection strategy for both the Core and Uncore components, tailored to each function in the workflow, based on the propagated latency up to the current stage. $\Omega$kypous uses a global gray-box model, to estimate with negligible error (7% MAPE) the latency and power consumption per configuration, to control the serverless chain propagated latency per stage, tuning the Core and Uncore frequency across the enriched solution space of configurations. $\Omega$kypous outperforms the related SotA in terms of QoS violations and power efficiency.
Contributions. Our main contributions are: i) Detailed analysis of the combined effects—latency and power consumption—of Core and Uncore frequency scaling on serverless computing functions, alongside a comparison with the CPU-quota scaling approach. ii) We design a global gray-box model for estimating function latency that considers the dynamicity of both the Core and Uncore parts of the processor. iii) We introduce $\Omega$kypous, the first Core and Uncore frequency controller for QoS-precise serverless workflows. iv) We provide an extensive comparative evaluation study w.r.t. the respective SotA DVFS frameworks, validating the practicality and the effectiveness of $\Omega$kypous, showing its impressive capabilities in nearly eliminating QoS violations while achieving up to 21% power savings. v) In-depth characterization of $\Omega$kypous, including decomposition, noise assessment, and responsiveness analysis.
2. Related Work
Table 1. Qualitative comparison between state-of-the-art and $\Omega$kypous
Pegasus(lo2014towards, ) Gemini(gemini2020zhou, ) ReTail(retail, ) SeqClock(seqclock, ) Aquatope(zhou2022aquatope, ) DVFaaS(tzenetopoulos2023dvfaas, ) $\Omega$kypous
Modeling - NN Classification Linear Regression - Bayesian Optimization - Interpolation+LR
Objective (s.t. QoS) Power Min. Power Min. Power Min. CPU util. Min. Cost Min. Power Min. Power Min.
Technique RAPL CFS CFS Core Scaling Core Scaling CFS CFS & UFS
Scope Service Service Service Workflow Workflow Workflow Workflow
$\mu$arch-aware
Propagation-aware
DVFS-enabled power management: Various research efforts have been made to optimize the power consumption of latency-critical applications using DVFS (retail, ; lo2014towards, ; kasture2015rubik, ; gemini2020zhou, ; adrenaline, ; tzenetopoulos2023dvfaas, ). Pegasus (lo2014towards, ) was one of the first works to use package-level power limit adjustments via RAPL for managing latency objectives in server systems, however, its coarse-grained approach and reliance on application-specific parameters limit its flexibility. In response, Rubik (kasture2015rubik, ) and Adrenaline (adrenaline, ) shifted focus to request-level granularity in power management. However, Adrenaline’s reliance on application-level features for identifying long requests and Rubik’s conservative approach focusing on the tail of latency distributions render them unsuitable for diverse workload scenarios. More recently, Gemini (gemini2020zhou, ) and Retail (retail, ) proposed CFS-driven strategies again at request-level. Gemini relies on black-box, application-specific neural networks for request latency prediction, which can introduce huge overheads to the milliseconds requirements of serverless. On the other hand, Retail relies on linear regression techniques, which, although simpler and more generalizable, may not fully capture the dynamic nature of complex cloud environments. Despite their respective modeling limitations (which we further discuss in §5.1), all these works operate at microservice- or request-level granularity, overlooking the advantages presented by serverless computing. Moreover, their focus remains on CFS-driven mechanisms, leaving the impact of UFS on power management relatively unexplored. DVFaaS(tzenetopoulos2023dvfaas, ) is most probably the closest work to ours, which leverages the intermittent nature of serverless workflows to address their end-to-end latency requirements while minimizing power consumption. However, DVFaaS still relies on CFS-driven power management, while also following a reactive strategy for adjusting core frequencies based on slack propagation.
QoS-aware strategies: Several works have also been proposed to optimize resource allocation for latency-critical applications under QoS constraints (chen2019parties, ; patel2020clite, ; sfakianakis2021skynet, ; nishtala2020twig, ; zhang2021sinan, ). While these works effectively address resource optimization to meet latency requirements, they do not focus on power management as a key optimization goal. Moreover, they operate at microservice-level, overseeing the opportunities offered by the serverless paradigm. Latest research efforts target explicitly serverless workflows, either considering the minimization of cost expenses (moghimi2023parrotfish, ; wen2022stepconf, ; eismann2021sizeless, ; zhou2022aquatope, ) or the regulation of the end-to-end latency (seqclock, ; stojkovic2023specfaas, ; llama, ; klimovic2018pocket, ; zhao2021understanding, ; da2019using, ). SeqClock(seqclock, ) and Aquatope(zhou2022aquatope, ) are closely related to our work. However, while they employ CPU-quota scaling to meet QoS requirements in serverless workflows, they do not prioritize power optimization. SeqClock follows a function-agnostic, step-wise, closed-loop approach to regulate end-to-end execution latency of serverless workflows, while Aquatope (zhou2022aquatope, ) employs Bayesian Optimization to build black-box models that minimize CPU time usage while addressing the end-to-end QoS constraints. However, as we later show in §3, DVFS strategies appear to be a more suitable configuration knob for power management compared to CPU quotas.
$\Omega$kypous value & differentiators: $\Omega$kypous introduces a practical DVFS approach that uses a single, lightweight global gray-box model to predict function execution latency, taking also into consideration the impact of frequency scaling on the Uncore parts (UFS) of the processor infrastructure. It adapts dynamically to any performance variability by utilizing closed-loop control for per-core DVFS, leveraging the by-design step-wise, fine-grained nature of serverless chains. It is designed for QoS precision, adapting to multiple levels of QoS requirements, while minimizing the power consumed. Table 1 summarizes the most closely related works, providing a qualitative comparison and the position of $\Omega$kypous with current state-of-the-art works.
3. DVFS strategies and Serverless: Open Challenges & Opportunities
Challenge 1: Vertical CPU scaling vs. Core DVFS: Between CPU- and voltage/frequency-scaling knobs, which is the more effective and robust method for regulating execution time while minimizing power consumption?
CPU-quota scaling has been used in recent studies for latency regulation of serverless workflows (seqclock, ; zhou2022aquatope, ). To highlight the potential of CPU quotas compared to DVFS strategies, we examine their impact on power consumption and execution latency. Specifically, we conduct a detailed experiment across various serverless functions, evaluating the average latency and power consumption over 100 invocations. For CPU quotas, we allocate 8 vCPUs per function and we vary the quota limit within the range of 0.5 to 8, under the default frequency scaling policy (ondemand Linux governor). For DVFS, we only vary the core frequency component (CFS), within the range of 1.2 to 2.5 GHz and we allocate two vCPUs per function [^1] Without loss of generality, for the rest of the paper, we consider two vCPUs available per function., which is the most common configuration for serverless deployments (wang2018peeking, ). The detailed experimental setup is described in §6.
Figure 2 shows the results, where we observe a notable trend in the comparison between CPU quotas and CFS. Interestingly, even though CPU quotas allow for more cores to be utilized, CFS reveals a similar range but with more fine-grained latency values while simultaneously achieving reduced power consumption. This suggests a more efficient trade-off between latency and power within the CFS approach. The observed increase in power consumption with larger CPU quotas is attributed to the operational characteristics of the ondemand governor, which dynamically scales the frequency up/down in response to changes in core utilization. Moreover, we also see that diverse applications exhibit substantial variations in their CPU quota requirements; for instance, thumbnailer demands as low as 0.5 CPU quota for achieving low-latency performance, while graph-pagerank requires as much as 2 CPU quota respectively. These disparities underscore the inherent challenges in achieving performance predictability and power proportionality.
[…]
Figure 2. Impact of CPU-quota and DVFS on the latency ($90^{th}$ percentile) and Power consumption of functions.
Challenge 2: Adequacy of existing frequency governors: Are the current DVFS policies sufficient for QoS-aware power management? One of the key components of power management in modern processors is the implementation of governors, which are software modules in the Linux kernel that control the frequency and voltage of the CPU based on system demand and power constraints. The four most common CPU governors are i) powersave, ii) performance, iii) ondemand, and iv) conservative available through the CPUFreq subsystem in Linux. To assess the impact of each governor on performance and power consumption, we sequentially invoke ten functions from (sebs, ) four separate times, each time using exclusively one of the aforementioned governors. Our findings indicate that powersave and performance governors anchor the CPU frequency at its minimum and maximum limits respectively. In contrast, the ondemand and conservative governors adjust the CPU frequency based on CPU utilization, with the latter showing a more gradual response to utilization changes, thereby promoting reduced power consumption. This analysis reveals that Linux governors behave identically regardless of any specified target QoS requirements, i.e., current DVFS governors in the Linux kernel are QoS-agnostic.
[…]
Challenge 3: Investigating functions’ performance sensitivity under Core- & Uncore-Frequency Scaling Next, we examine the impact of CFS and UFS on the latency and power consumption of serverless functions. Specifically, we deploy the thumbnailer and video processing functions from (sebs, ) and measure their tail latency and power consumption for various configurations of the Core and Uncore frequencies available in our system (as described later in §5.5). Figure 12 summarizes the results, where the green-shaded lines indicate the configured Uncore frequency level. The red line represents the system-managed UFS (default configuration), which utilizes Intel®s firmware to vary the Uncore frequency dynamically according to the CPU usage (hill2010uncore, ; gholkar2019uncore, ). Based on these outcomes, we answer two fundamental questions:
(a) Impact on latency
(b) Impact on power
Figure 3. Impact of Core (x-axis) and Uncore frequency on tail latency and power consumption.
1) Is direct linear proportionality between Core frequency and latency a reasonably safe assumption? Prior research (kasture2015rubik, ; gemini2020zhou, ; chou2019mudpm, ) assume linear proportionality between application latency and applied core frequency. For example, Gemini (gemini2020zhou, ) calculates latency values at different frequency levels by assuming that the product of latency and frequency remains constant. However, as demonstrated in Fig. 3a the relationship between CFS and latency varies per function. For example, in the case of video-processing, scaling the Core frequency from 1.2GHz to 2.5GHz results in $\approx$ 50% tail latency reduction while in thumbnailer it is only improved by 30%. Therefore, given the considerable variability in the frequency-to-latency relationship across different serverless functions, relying on a linear (one-to-one) proportionality is not a universally applicable assumption. However, as depicted in Fig. 1, UFS allows for leveraging iso-latency configurations for lower power consumption. Notably, in the 2.5GHz core frequency scenario, when explicitly applying UFS, power consumption is reduced by 23.5%, with the latency degrading by only 0.3%, 10% for the thumbnailer) and video-processing respectively.
[…]
Challenge 4: Functions’ latency modeling under DVFS
Recently, Azure released a trace dataset (zhang2021faster, ) from its datacenter, comprising invocations of 424 distinct functions recorded over a two-week period. As illustrated in Figure 4,
Figure 4. Function invoc. count CDF from Azure Trace (zhang2021faster, ).
half of these functions were invoked less than 30 times during the provided time period, highlighting the existence of “small-data” scenarios in serverless. Given the variability in input sizes within serverless computing (romero2021faa, ), models must be adept at predicting latency for diverse configurations. This observation pinpoints the need for latency estimation models that are accurate yet lightweight, offering efficient retraining capabilities and low-overhead inference suitable for such “small-data” contexts. Unfortunately, in the context of predictive methodologies this is not yet the case, e.g., Retail (retail, ) requires at least one execution for each function and each available frequency configuration (under fixed input sizes).
[…]
4. $\Omega$kypous: Problem formulation
$\Omega$kypous Rationale: $\Omega$kypous relies on the intermittent nature of serverless workflows, with its core idea being to leverage slack propagation effects between functions. This approach enables the system to explore optimal CFS/UFS configurations for each function, ensuring power consumption is minimized while closely satisfying target latency QoS constraints, i.e., being QoS-precise. Figure 5 shows a simplified motivational example, which considers a serverless chain consisting of 5 functions and a conditional branch. In this case, the target QoS corresponds to the worst-case scenario, i.e., the cumulative latencies of each function that forms part of the critical path. However, during actual execution, the observed latency can deviate, leading to negative and positive timing slacks, which then propagate through the entire workflow. Negative slacks, indicating delays, can be a result of noise effects e.g. interference (chen2019parties, ; zhao2021understanding, ), or inaccuracies in latency predictions (zhou2022aquatope, ; gemini2020zhou, ). Positive slacks, which imply time saved, may result from similar latency prediction errors or system noise. The intermittent nature of serverless workflows facilitates the identification of timing slacks, which can be strategically utilized to apply “just-enough” frequency to each function. This approach can optimize power consumption while closely adhering to target QoS constraints (QoS-precision).
Power efficiency subject to QoS: Power consumption is a critical concern for modern cloud infrastructures, with power capping being a first-class strategy followed in high-end data centers (kontorinis2012managing, ; lo2014towards, ). $\Omega$kypous optimizes power consumption of serverless workflows under QoS constraints by targeting the following problem specification:
[…]
Specifically, for a workflow comprising of $n$ functions, we aim to minimize the power consumed per function $P_{k}$, while preserving the end-to-end QoS latency requirement or latency target – $L(f_{1},..f_{k},..f_{n})\leq T$.
Pre-deployment Latency Budgeting: To project this high-level objective to our problem, we need first to employ a latency budgeting model, i.e., a mechanism to distribute the workflow’s end-to-end latency budget to several per-function latency constraints. As shown in the top of Figure 7, a serverless workflow can be represented as a DAG of functions. DAG’s execution is dynamic, thus most of the time the execute paths are defined during runtime, based on the output of the intermediate functions. Nevertheless, a DAG can be decomposed in several mutually exclusive paths, as shown in the same Figure (left part of pre-deployment phase). To construct the latency budgeting model, given the end-to-end latency constraint $T$, we need first to consider every possible serverless chain or path $i$. We calculate the critical path (CP) $\max_{i}{\sum{\min{l_{k}^{i}}}}$, based on the maximum end-to-end execution time of chains, considering the sum of the minimum latency $l_{k}$ that can be achieved per function $k$. Given the various DVFS strategies of each function k we can select any combination of target function latency $t_{k}$ that i) meets the end-to-end requirement $T$ ($\geq CP$), and minimizes the slack, or deviation from the target $T$:
[…]
This procedure allows us to define the QoS target latency $t_{k}^{cp}$ per function $k$ that resembles the chain of the critical path (right part of pre-deployment in Figure 7).
Figure 5. Slack propagation management for QoS-precise power-efficient workflow execution.
Dynamic Latency Budgeting: Based on the DAG representation and the established per-function latency budgeting, $\Omega$kypous dynamically adjusts the budgeted latency per function at runtime, taking into account slacks propagated from previous functions. To avoid speculation and the corresponding misprediction costs (stojkovic2023specfaas, ), we adopt a less aggressive strategy (see §5.2) that tries to meet the accumulated QoS requirement up to the current stage of the workflow. For example, in the DAG of Figure 5, after function $f_{2}$ returns and $f_{4}$ is determined as the next function to be invoked after the conditional branch, we have to select a configuration for the
Figure 6. Fan-Out / Fan-In scenario.
execution of function $f_{4}$ that will meet the following objectives: i) remain under the QoS constraint at stage $4$ and ii) minimize the power consumed. We note that the adopted dynamic time budgeting approach methodology remains consistent and applies also to the Fan-out/Fan-in serverless workflow scenario. For instance, considering the workflow in Fig. 6, we choose a configuration that effectively mitigates the accumulated slack for each one of the concurrent functions $f^{1}_{1-n}$, while minimizing power consumption.
5. $\Omega$kypous Design
$\Omega$kypous addresses the problem of QoS-precise power management for serverless workflows. We design $\Omega$kypous based on the opportunities identified in §3, by leveraging advanced DVFS, i.e., cooperative CFS and UFS, as the knob to drive optimization. Figure 7 shows an overview of $\Omega$kypous design, which we analyze in the remainder of this section.
5.1. Latency & Power modeling
Serverless functions exhibit significant dynamicity in their configurations, both in terms of hyperparameters (e.g., compression rates for video processing) as well as in the size of inputs, which can vary greatly across different invocations. Given this variability, an effective latency estimation model must be capable of capturing such parameters to make accurate predictions. To this end, $\Omega$kypous employs a two-staged, lightweight modeling approach, that leverages hardware performance monitoring counters (PMCs), a practice commonly used in data centers (wang2022characterizing, ), which provides a prominent way to characterize system performance and guide decision-making (bianchini2020toward, ; masouros2023adrias, ). The first step maps the impact of varying payload sizes to these performance counters. The gathered data are then utilized by a grey-box model, which integrates mechanistic insights with theoretical principles, to predict the latency and power consumption of serverless functions under various Core and Uncore frequency settings. Opposed to previous approaches that often require a distinct model for each function and frequency configuration (retail, ), or rely on a direct proportionality assumption between frequency and latency (gemini2020zhou, ), $\Omega$kypous employs a global modeling strategy. Specifically, driven by Opportunity 4, we construct a single model that is applicable across all functions, enhancing its efficiency and scalability in diverse serverless environments.
5.1.1. Payload to PMCs interpolation:
The performance of functions is highly affected by variations in the size of input data. The data size can reflect the number of nodes in graph processing or the byte size in image processing. To identify which input knob influences function latency, a preliminary correlation analysis is required (retail, ). Indicatively, Figure 8a shows the measured execution time of graph-bfs, dynamic-html and dna-viz functions under the selected scaled payload sizes. As depicted, the execution time of these functions exhibits a linear scaling behavior relative to the payload size, with the execution variability increasing for higher payload sizes. Still, representing payload size in a global modeling approach can be challenging. Due to the diversity of input types across functions – both in terms of size and content – it becomes impractical to encode payload size into a unified scale. Moreover, categorizing payload sizes, such as mapping sizes to discrete labels (e.g., notate datasets under 1MB as “small”), not only restricts the model to a discrete space but also fails to provide flexibility for interpolating to unseen payloads.
Figure 7. $\Omega$kypous overview.
To address this challenge, we employ a strategy where different payload sizes are mapped to their corresponding PMCs under a default Core/Uncore frequency configuration [^2] For simplicity and without loss of generality, we set the default frequency to 1.4GHz for both Core and Uncore components.. These PMC mappings are then utilized by our grey-box model (refer to Section 5.1.2) to make predictions for other frequencies. PMCs allow to have unified and function-agnostic interpretation of payload sizes, while they closely correlate with both payload size and execution time of functions. Figure 8c shows the Pearson correlation (cohen2009pearson, ) of all the collected PMC with the payload size and execution time of each function. As shown, despite the different functionality and runtime behavior of the three examined functions, there is a set of important PMCs, e.g. Figure 8c , that exhibit high correlations (¿0.7) w.r.t. payload size, thus allowing PMCs to characterize the payload-related latency values.
Specifically, $\Omega$kypous maintains a meta-data table that records the PMCs from previous executions of each function, where each entry is associated with its corresponding payload size. In instances where $\Omega$kypous encounters a new request with a payload size not previously observed, the system estimates its PMCs by performing a linear interpolation based on the nearest data points available in the meta-data table. Figure 8c evaluates the interpolation’s achieved accuracy over the overall function set of $\Omega$kypous. As shown, the proposed interpolation scheme manages to deliver high accuracy over PMCs of high and moderate predictive importance [^3] We characterize PMCs’ importance by the amplitude of the weights derived after training the linear interpolation model. with MAPE ranging between 0.01-0.05. Last, $\Omega$kypous considers functions with different hyperparameters (e.g., compression rate in compression function) as separate entries in the meta-data table, i.e., different functions.
5.1.2. Grey-box modeling of latency and power
Figure 8. a) Impact of different payload sizes on the execution time of serverless functions. b) Pearson correlation of all the examined hardware performance counters with the payload size of each function. c) Interpolation error (MAPE) for the most important PMCs, as designated by the LR regressor (§5.1.2).
After estimating the PMCs for the invoked function we use them as a means to predict the execution latency [^4] A similar approach is also followed for power modeling. under different CF, and UF configurations. We employ a single grey-box model that formulates execution latency as a function of the Core $F^{c}$, Uncore $F^{u}$, and DRAM frequencies $F^{d}$, and total number of cycles $W^{c}$, $W^{u}$, $W^{d}$ concluded on Core, Uncore, and DRAM parts. Prior research  (choi2004fine, ; deng2012coscale, ; sarood2014maximizing, ) has shown that execution latency depends on the workload distributed across the CPU, the on-chip memory, and the off-chip DRAM memory components, thus it can be modeled as:
(3) $L(F^{c},F^{u})=\frac{W^{c}}{F^{c}}+\frac{W^{u}}{F^{u}}+\frac{W^{d}}{F^{d}}$
However, accurately counting the Core, Uncore and DRAM working cycles on a superscalar architecture is a quite complex procedure. To this end, we transform Equation 3 to represent $W^{c}$, $W^{u}$, $W^{d}$ as the sum of relevant PMCs over time [^5] Due to our hardware (§6) not supporting DRAM frequency scaling, we treat its impact as constant value in our design.. Specifically, for each function, we map $W^{c}$, $W^{u}$, $W^{d}$ to the PMCs of the default CF,UF frequency ($F_{base}^{c},F_{base}^{u}$) configuration (§5.1.1). Based on this, Equation 3 is transformed to the following:
(4) $L(F^{c},F^{u})=\frac{\sum_{i=1}^{N}{c_{i}\cdot P^{c}_{i}}}{F^{c}}+\frac{\sum_{i=1}^{N}{u_{i}\cdot P^{u}_{i}}}{F^{u}}+\sum_{i=1}^{N}{d_{i}\cdot P^{d}_{i}}$
In our model, $P^{c}_{i},P^{u}_{i},P^{d}_{i}$ represent the PMCs chosen to reflect the working cycles of the Core, Uncore, and DRAM components, respectively. The weights of each PMC in modeling the working cycles ($W^{c}$, $W^{u}$, $W^{d}$) are denoted by $c_{i},u_{i},d_{i}$. We carefully select the most relevant PMCs for each component based on correlation analysis and insightful information for current literature (chen2023joss, ). To determine the coefficients $c_{i},u_{i},d_{i}$, we employ regression analysis. Specifically, we examine various regression methods including Linear Regression, Decision Trees, and Extreme Gradient Boosting (xgBoost). We asses each method by evaluating the end-to-end case, i.e., we test each models with function/payload pairs not encountered during training. The PMCs for these payloads are calculated using interpolation (§5.1.1) and are propagated to our grey box model to estimate the final latency. Table 2 shows the respective results per regression approach and per examined function (§6), by reporting the Mean Absolute Percentage Error (MAPE). As shown, there is no clear winner among the examined cases. However, Linear Regression present the more robust MAPE across all cases and also provides the lowest error in the majority of the cases.
Table 2. End-to-End MAPE (%) for different regressors.
vp dh cmp ir gb gp gm th dv
LR 1.8 14.8 9.6 3.7 9.3 2.8 4.9 5.5 3.3
DT 0.7 19.2 16.5 0.8 16.7 2.9 14.7 7.4 15.9
XGB 0.8 13.4 14.8 1.4 12.7 4.5 6.1 9 15.3
5.2. Adaptive controller
$\Omega$kypoys targets serverless DAGs, which are composed of multiple functions. Each function is deployed and serves users’ requests without reporting any application-level metrics related to performance, during the execution. Thus, to avoid function-specific configurations for checkpointing, and extra overhead induction, the reporting interval is equal to the function serving time. Exploiting this opportunity, we employ a closed-loop controller, that is activated between function invocations. Let $t_{k}$ be the QoS-target latency and $l_{k}$ be the actual measured latency, for the function in stage $k$ of the serverless chain. In Eq.5 we define the cumulative (up to stage $k$) target $T_{k}$, measured $L_{k}$ latency as well as the violation time or slack $V_{k}$.
(5) $T_{k}=\sum_{i=1}^{k}{t_{i}},\quad L_{k}=\sum_{i=1}^{k}{l_{i}},\quad V_{k}=T_{k}-L_{k}$
Considering the target control loop, $L$ forms the process variable, $T$ is the setpoint, and $V$ forms the error. We design our controller according to two principles: i) rigorously restrain the system from any cumulative QoS-target violation ($V(k)\leq 0$), and ii) minimize the area between measured and target QoS, i.e., the unexploited latency slack overtime ($\min{\sum_{i=1}^{k}{V(i)\times l(i)}}$). The output of the controller is $u(k)=K_{p}V(k-1)$, where $K_{p}$ is the proportional gain, which is adjusted dynamically as a function of $V(k-1)$ according to Eq. 6.
(6) $K_{p}=\left\{\begin{array}[]{ll}R&\mbox{if }V(k-1)>0\\ r_{1}V(k-1)^{2}+r_{2}V(k-1)+r_{3}&\mbox{if }V(k-1)\leq 0\end{array}\right.$
where $R$ is a constant to overcome any violation as soon as possible, while $r_{1},r_{2},r_{3}$ are parameters, extracted through a grid search approach, that construct the quadratic function.
5.3. Conflict Resolver
The Conflict Resolver component is designed to efficiently manage DVFS conflicts arising from differing frequency requirements among concurrently running functions. Conflicts most commonly occur within the shared Uncore components on a per-socket basis but can also emerge in scenarios where different functions are co-located on the same physical cores. To manage these conflicts, the Conflict Resolver employs two priority queues: one for Core frequency requests (CFS) and another for Uncore frequency requests (UFS). In these queues, priorities are assigned based on frequency values, with higher frequencies receiving higher priority. When the Adaptive Controller submits a new CF/UF configuration, the Conflict Resolver adds these frequency values to the corresponding priority queues. It then checks whether the highest-priority frequency in the queue is lower than any frequency currently requested by the running functions. If so, the existing frequencies are maintained to avoid disrupting the ongoing operations of these functions (Fig. LABEL:fig:conflict-resolver). However, if a new, higher-priority frequency request appears in the queue, suggesting no conflict with running functions, the Conflict Resolver then signals for these new frequencies to be applied to the system (Fig. LABEL:fig:conflict-resolver). If the queues are empty, the lowest frequencies are applied. This hierarchical approach ensures that frequency changes are appropriately adjusted, respecting the operational requirements of all active functions.
5.4. Discussion
$\diamond$ How does $\Omega$kypous deal with performance variability errors due to mispredictions, performance variability and/or queuing effects? Except for latency mispredictions, applications co-location can introduce resource interference, potentially leading to performance variability. Consequently, the additional latency overhead denoted as $E$ arising from this variability has the potential to adversely affect the overall end-to-end latency of the serverless workflow. $\Omega$kypous, by employing the adaptive controller, can reactively mitigate any latency overhead, taking it into account in the slack calculation: $V(k)=T_{k}-(L_{k})+E(k))=>V^{\prime}(k)=V(k)+E(k)=T(k)-L(k)$.
Figure 9. Serverless functions $95^{th}$ percentile latency on varying request per second (rps) and concurrency levels.
$\diamond$ How does $\Omega$kypous handle cold-start effects? $\Omega$kypous does not provide a component that directly manages cold-starts. Nevertheless, the Azure trace reveals that the execution time of functions deployed on their premises greatly varies from seconds to minutes (zhang2021faster, ; shahrad2020serverless, ). In cases of workflows comprising long-running functions, cold-starts can be eliminated through slack management, as described in §4. On the contrary, in cases of serverless chains containing short running functions, $\Omega$kypous can be integrated with frameworks targeted on mitigating cold start latency (zhou2022aquatope, ; roy2022icebreaker, ; fuerst2021faascache, ) by proactively pre-warming functions. These approaches address an orthogonal problem to ours, primarily focusing on minimizing idle resource usage.
5.5. $\Omega$kypous Implementation
We implemented $\Omega$kypous as an extension of Apache Openwhisk (openwhisk-site, ) over a Kubernetes (kubernetes, ) resource orchestrator.
Storage: We utilize MinIO (minio, ) object storage to store inputs/outputs of functions. MinIO is deployed on a remote node, to account for network overheads for data fetching (klimovic2018pocket, ).
Core/Uncore DVFS: We deploy one agent on each physical server, which receives requests and applies the corresponding Core and Uncore frequencies. Each agent exposes an HTTP endpoint by utilizing the Flask (flask, ) web framework, which can be consumed by the Conflict Resolver to alter the respective frequency configuration. Our Intel® server (§6), offers userspace per-core scaling configuration, i.e., per-core DVFS. To specify the desired core frequency explicitly, we disable the default scaling driver intel_pstate and utilize the acpi_cpufreq. We select the userspace as the scaling governor and change the frequency per core through the sysfs entry /sys/devices/system/cpu/cpuX/cpufreq/scaling_setspeed.
[…]
Core Pinning: We ensure a one-to-one mapping between serverless functions and Core frequencies by allocating specific vCPUs for each function, with each vCPU mapped to a physical core on the server. This process is managed through a Pub/Sub system using RabbitMQ (rabbitmq, ). Function-to-core mappings are published during function deployment and consumed by agents on the virtual hosts, which pin functions to designated cores using Docker’s cpuset-cpus flag.
Latency, Power & PMC Monitoring: The latency per function is inherently measured in the core logic of $\Omega$kypous between consequent invocations. Moreover, power consumption is monitored through the RAPL-based (david2010rapl, ) Intel® Performance Counter Monitor(pcmgit, ) in $20ms$ intervals. Last, for PMCs collection, we deploy one agent per physical host, which is triggered through a separate endpoint on the physical servers, and monitors the performance events of each core on the system using linux-perf.
6. Experimental Setup
(a) Barycenter Analysis Derived from K-Means Clustering of 60-Second Segments in Azure Trace Data (zhang2021faster, ).
(b) Exploring the performance of Linux frequency governors and state-of-the-art power managers, alongside $\Omega$kypous, across various clusters. The figure showcases distributions of QoS Precision (top) and Power (bottom) observed during invocations of serverless workflows within each examined cluster.
Figure 10. Comparative analysis of $\Omega$kypous with other power management governors for different invocation patterns, as derived from the Azure Trace dataset (zhang2021faster, ).
Examined Infrastructure: We deploy four VMs with 10 vCPUs and 16GB each, on top of two Intel® Xeon® E5-2658A servers, a common setup found in public serverless solutions (wang2018peeking, ). We pin each VM on a separate socket to allow for granular Uncore frequency, as described in §5.5. We also disable Hyper-Threading, as CFS is applied per physical core.
Target Benchmarks: We employ 8 different functions from the Serverless-Benchmark-Suite (sebs, ), which represent typical serverless workloads and align with previous studies (zhang2021faster, ; shahrad2020serverless, ). Specifically, we utilize the following functions: video-processing (vp), dynamic-html (dh), compression (cmp), img-class (ic), graph-bfs (gb), graph-pagerank (gp), graph-mst (gm), thumbnailer (th) and dna-visualization (dv). We configured a vCPU limit of 2 cores per function, in line with the median configuration reported in (joosen2023does, ). This setting also enables us to include scenarios of core co-location due to the imposed vCPU cap. Moreover, based on these functions, we generate different workflows by building different permutations to simulate realistic workflow variability, including chain, fan-out/in, and mixed compositions. The lengths of these workflows are modeled based on distribution patterns reported for applications running on Azure (shahrad2020serverless, ).
Workflow QoS calculation: To establish the target QoS for each workflow, our initial step is to determine the maximum input load that each function can sustain. This is achieved by starting with a low request-per-second (RPS) rate and gradually increasing the request rates for each function across various levels of concurrency using the ondemand frequency governor. We continue this process until reaching a saturation point. This point is characterized by an “explosion” in tail latency indicating that the RPS has exceeded the system’s capacity to handle requests effectively. Figure 9 presents the results for each function under varying RPS and concurrency levels. Each cell represents the $95^{th}$ percentile latency for its corresponding RPS and concurrency level, and cells with blue borders highlight the maximum RPS that each function can handle. To define the QoS for each workflow, we calculate the cumulative $95^{th}$ percentile latency of all functions comprising the workflow. Additionally, the maximum load that the workflow can sustain is determined by identifying the lowest RPS among the functions within it.
Load generation and deployment approach: We generate load to our system by utilizing invocation traces sourced from Azure (zhang2021faster, ). These traces, available in the latest Azure Public Dataset, are labeled according to their associated applications. We define an application invocation as the initial invocation of the first function within each application. In our evaluation of $\Omega$kypous, we opted for a methodology distinct from downsampling, which would entail sacrificing trace granularity at a finer time scale (joosen2023does, ). We employed time series K-means clustering on individual segments, each spanning 60 seconds to create ten clusters. This approach allows us to preserve the dynamicity of the traces while facilitating the identification of key patterns and behaviors(joosen2023does, ). We omit the clusters indicating inactivity (cluster 0), and the ones that are found only on single-stage applications (cluster 7). Figure 10a shows the barycenters of each cluster, showing the different invocation patterns identified through the clustering approach. Last, we down-scale the invocation rate, to be consistent with the maximum RPS described in Figure 9. To generate invocation requests, we sample segments from each of the eight remaining clusters, to construct an experimental scenario of 160 minutes. For each selected sample, we use serverless workflows of the same length as the one reported in the trace, using random permutations of serverless functions, as described above.
7. Evaluation
7.1. $\Omega$kypous vs power management schemes
In this section, we quantitatively compare $\Omega$kypous efficiency w.r.t. other power management schemes. Figure 10b presents aggregated results of our experiment regarding to each examined cluster of the Azure Trace. The top row of Figure 10bb shows the distributions of the tail latency ($95^{t}h$ percentile) normalized to the QoS target for each workflow. In the second row, we plot the distributions of power consumption during the execution of the sampled workflows for each cluster. In the bottom of Fig. 10b, we depict the $90^{th}$ percentile of power consumption.
Typical Power Management Schemes: We assess two Linux frequency governors, ondemand and conservative, excluding performance and powersave for their static frequency settings, leading to high power consumption or QoS violations. Ondemand, the default governor, exhibits inefficient power usage, while conservative, with a less aggressive frequency fluctuation, shows improved consumption. significantly enhances power efficiency, reducing consumption by an average of 16% and 11% compared to ondemand and conservative, respectively. This improvement is more pronounced in scenarios with lenient QoS requirements or the omitted (almost inactive) cluster trace.
State-of-the-art power managers: Next, we compare $\Omega$kypous with the strategies of two SotA CFS-based power managers. Gemini (gemini2020zhou, ) and ReTail (retail, ) are designed to manage distinct micro-services. To transform their implementation to the serverless workflow setup we consider two different approaches. First, we implement them considering the entire serverless workflow as a single service (ReTail, Gemini), which is managed as a whole. Second, we implement them in a decomposed manner, in which each serverless function is managed independently by a separate controller (ReTail+, Gemini+). For the latency estimations, we use the average values of the measured execution times (for each frequency level for ReTail, and the base frequency for Gemini).
QoS-violations: As depicted in Fig. 10b, Gemini and Retail exhibit QoS-precise end-to-end latency in certain clusters (3 and 9) with the chosen frequency. However, performance variability arises from the diverse functions and I/O inter-communication overheads, leading to performance variability that cannot be anticipated at the initiation of the workflow. Gemini presents the most violations, occurred from the inaccurate methodology (assuming linear proportionality).
Power consumption: $\Omega$kypous improves power consumption up to 17% and 18% (9% and 16% on average) compared to Retail and Gemini respectively, as shown in Fig. 10b The factors that lead to this result include: i) The more fine-grained power management enabled by application decomposition into functions. ii) The introduction of UFS management which furtherly reduces power consumption as discussed in §7.2. Notably, as shown in the bottom of Fig. 10b, the $90^{th}$ percentile of power, is reduced up to 21% compared to Retail. This would allow for cluster-wide power-capping (RAPL) strategies without increasing the QoS violations.
7.2. $\Omega$kypous characterization
Responsiveness to change: To further explore the operational dynamics of $\Omega$kypous, particularly in terms of its ability to manage Core and Uncore frequencies under different conditions, we have developed a specific serverless workflow. We execute this workflow under varying request-per-second (RPS) levels and tested against three distinct Quality of Service (QoS) levels. Figure 11 illustrates these experiments, showing the tail latency corresponding to each QoS constraint (indicated by red lines) along with the fluctuations in Core and Uncore frequencies over time as RPS increases. The results demonstrate that the $95^{\text{th}}$ percentile latency consistently stays below the pre-set QoS thresholds, showing $\Omega$kypous’s effectiveness in satisfying QoS requirements through its DVFS strategies. Notably, we observe a correlation between higher frequency values and stricter QoS requirements. Specifically, both Core and Uncore frequency exhibit an upward trends in their distributions as the RPS intensifies and the QoS targets become more demanding. Particularly under the strictest QoS constraint (right figure), we also observe instances where $\Omega$kypous maintains low Core frequencies, highlighting the significant role of Uncore Frequency in latency regulation.
Impact of UF management: At next, we evaluate the additive impact of UFS management. We implement a CFS-only version of $\Omega$kypous ($\Omega$kypous-CFS), which only manages the Core frequency. In Fig. 12a, based on the Azure Trace clusters (Fig. 10a), we compare the power distributions of the two solutions across all clusters. As shown, we separate the power consumption on DRAM and PROC components. The spikes in power (in time windows equal to each workflow’s execution time) are further decreased by $\approx 15\%$. The average value is decreased by 5%/3% for DRAM, and PROC. With $\Omega$kypous-CFS, similar to $\Omega$kypous, addressing the QoS requirements, the application of strategies such as power capping in both components (DRAM, PROC) would be a better fit for the UFS-enabled version.
Resilience to execution noise: We further evaluate the $\Omega$kypous controller’s ability to mitigate noise injected during the workflow execution. Figure 12b, illustrates the QoS latency violations on different noise levels (x-axis), applied on different fractions of the workflows (y-axis), on two different workflow types that could affect the result (right/left skewed). The left-skewed workflow, which concentrates the relatively larger latencies at the end of the workflow can adhere to the QoS requirement almost for every noise level when applied to half of the functions. On the other hand, for the right-skewed workflows, any error is harder to compensate since $\Omega$kypous does not have available many control steps to absorb the injected noise. In such cases, strategies such as dynamic QoS target adaptation (retail, ) or PMC resampling are required.
Figure 11. $\Omega$kypous Responsiveness.
(a) CFS-only vs. CFS/UFS
(b) Noise tolerance
Figure 12. Characterization.
[…]
8. Conclusion
In this paper, we introduced $\Omega$kypous, a novel QoS-precise DVFS framework for serverless workflows. $\Omega$kypous addresses the growing prominence of latency propagation effects in serverless architectures by employing closed-loop control to regulate end-to-end latency while minimizing power consumption through the optimized adjustment of Core and Uncore frequency combinations. A grey-box modeling approach is introduced to accurately predict execution latency and guide the coordinated CFS/UFS control decisions. We thoroughly evaluate $\Omega$kypous over the recently released Azure Trace data and we showed that $\Omega$kypous consistently outperforms state-of-the-art power managers, achieving up to 21% improved power usage on average compared to state-of-the-art power management schemes, while minimizing QoS violations and being resilient to external noise injection.